Given a metric d on a set X prove that there exists an equivalent bounded metric d' on X


1
Expert's answer
2021-07-15T05:32:50-0400

Solution;

Let d(x,y)=0

Let a,b and c be non negative numbers such that,

1+b+c"\leq" (1+b)(1+c)

"\frac{2+b+c}{(1+b)(1+c)}" "\leq" "\frac{2+b+c}{1+b+c}"

That is;

"\frac{1}{1+b}" +"\frac{1}{1+c}" "\leq" 1+"\frac{1}{1+b+c}" "\leq" 1+"\frac{1}{1+a}"

If 0"\leq{a}\leq{b+c}"

Add 1 to both sides of the above equation and rearrange;

"\frac{a}{1+a}" "\leq" "\frac{b}{1+b}" +"\frac{c}{1+c}" .......(i)

For any non negative numbers such that a"\leq"b+c

For x ,y ,z"\epsilon" X the none negative numbers a=d(x,y),b=d(x,z) and c=d(z,y) satisfy a"\leq" b+c by the triangle of inequality of metric d.

Equation (i) gives the triangle of inequality of d'(x,y)="\frac{d(x,y)}{1+d(x,y)}" in which d' is metric on X.